Are there multiple ways to convert a mixed number to a decimal?
I can only think of one way...
To convert a mixed number to a decimal, you would have to convert the mixed number to an improper fraction, then convert the improper fraction to a decimal

Found 3 solutions by Earlsdon, RAY100, rfer:
Answer by Earlsdon(6103)

You have identified one way but you don't really need to do anything at all with the integer part of a mixed number when converting to a decimal.
For example:
Convert to a decimal:
7 5/9 = 7+ 5/9 = 7.555... You just bring the integer part over and convert the fraction part to its decimal equivalent.
6 4/5 = 6+4/5 = 6.8

Answer by RAY100(1637)

Your method is correct. It does often result in larger numbers. But, not a problem with a calculator.
.
You might also realize that the whole number part of the mixed number can be considered part of an addition process. 1 1/2 is 1+1/2.
This allows us to convert the fractional part to a decimal, and then add to the whole number part, ie, 1+1/2 = 1+.5 = 1.5
.
Good luck and keep up the good work
.

Answer by rfer(10432)

In converting a mixed number to a decimal it is not necessary to convert to an improper fraction.
Just convert the fraction to a decimal equivalent such as shown below.
Divide the numerator by the denominator, properly place the decimal and show the whole number to the left of the decimal.
The whole number is the part of a decimal number that is to the left of the decimal such as in money.
1 1/3=1.33
12 1/4=12.25
1/200=.005